# X-MoGen: Unified Motion Generation across Humans and Animals

### v1

@@ 4.2 Qualitative Results
To further assess the effectiveness of our model, we qualitatively compare X-MoGen with three strong baselines: MoMask, MARDM, and AniMo, selected based on their quantitative performance. Figure 4 presents two representative examples. In the human motion case, X-MoGen accurately captures the prompt containing “running” and “cartwheel”, generating a coherent sequence with a smooth transition into the cartwheel. In contrast, MoMask reverses the intended action order, while MARDM and AniMo produce only the cartwheel, omitting the preceding context. Notably, the virtual tail nodes introduced for topological consistency do not degrade motion realism across all methods. In the unseen species case, X-MoGen generates motion that is both morphologically consistent and semantically faithful to the prompt. MoMask fails by abruptly morphing the animal into a human, MARDM misinterprets the prompt and produces unstable body rotations, and AniMo incorrectly generates a left-turn motion. These results demonstrate that X-MoGen effectively unifies human and animal motion synthesis, exhibiting strong generalization across species. Additional qualitative examples are provided in the supplementary material.
@@ 4.4 Applications
X-MoGen presents a unified framework for generating both human and animal motions from text within a shared representation space. The framework supports humans and 114 animal species, leveraging CGAE-generated, species-specific T-poses to generalize to new quadrupeds without additional training. This enables intuitive AI-driven motion creation and supports diverse applications including content production, biology-inspired simulation, digital storytelling, and interactive media (Abootorabi et al. 2025).
@@
X-MoGen enables smooth motion transitions between species. Given two motion sequences (e.g., a tiger walking and a human walking), the AE encoder extracts latent features, and mask tokens [M] are inserted at the transition point. M-Trans, with the diffusion head and AE decoder, generates realistic cross-species transitions. As shown in Figure 1, this enables smooth morphing between animal and human motions, supporting seamless cross-species motion editing and completion, valuable for animation and automated shapeshifting in film (Akber et al. 2023; Tessler et al. 2024).

### v2

@@ 4.2 Qualitative Results
To further assess the effectiveness of our model, we qualitatively compare X-MoGen with three strong baselines: MoMask, MARDM, and AniMo, selected based on their quantitative performance. Figure 4 presents two representative examples. In the human motion case, X-MoGen accurately captures the prompt containing “running” and “cartwheel”, generating a coherent sequence with a smooth transition into the cartwheel. In contrast, MoMask reverses the intended action order, while MARDM and AniMo produce only the cartwheel, omitting the preceding context. Notably, the virtual tail nodes introduced for topological consistency do not degrade motion realism across all methods. In the unseen species case, X-MoGen generates motion that is both morphologically consistent and semantically faithful to the prompt. MoMask fails by abruptly morphing the animal into a human, MARDM misinterprets the prompt and produces unstable body rotations, and AniMo incorrectly generates a left-turn motion. These results demonstrate that X-MoGen effectively unifies human and animal motion synthesis, exhibiting strong generalization across species.
@@ 4.4 Applications
X-MoGen presents a unified framework for generating both human and animal motions from text within a shared representation space. The framework supports humans and 114 animal species, leveraging CGAE-generated, species-specific T-poses to generalize to new quadrupeds without additional training. This enables intuitive AI-driven motion creation and supports diverse applications including content production, biology-inspired simulation, digital storytelling, and interactive media (abootorabi2025generative).
@@
X-MoGen enables smooth motion transitions between species. Given two motion sequences (e.g., a tiger walking and a human walking), the AE encoder extracts latent features, and mask tokens [M] are inserted at the transition point. M-Trans, with the diffusion head and AE decoder, generates realistic cross-species transitions. As shown in Figure 1, this enables smooth morphing between animal and human motions, supporting seamless cross-species motion editing and completion, valuable for animation and automated shapeshifting in film (akber2023deep; tessler2024maskedmimic).
@@ 5 Conclusion
Acknowledgments
This work was supported by the Zhejiang Provincial Natural Science Foundation of China (No. LZ24F030005), the National Natural Science Foundation of China (No. 62576308), the Fundamental Research Funds for the Central Universities (No. 226-2025-00167), and the Scientific Research Foundation of Sichuan Provincial Department of Science and Technology, China (No. 2024YFHZ0001).